### v1

@@ Abstract
Satisfying safety constraints is a priority concern when solving optimal control problems (OCPs). Due to the existence of infeasibility phenomenon, where a constraint-satisfying solution cannot be found, it is necessary to identify a feasible region before implementing a policy. Existing feasibility theories built for model predictive control (MPC) only consider the feasibility of optimal policy, since MPC itself can be viewed as an optimal controller that solves an optimal action at every step. However, reinforcement learning (RL), as another important control method, solves the optimal policy in an iterative manner, which comes with a series of non-optimal intermediate policies. Feasibility analysis of these non-optimal policies is also necessary for iteratively improving constraint satisfaction; but that is not available under existing MPC feasibility theories. This paper proposes a feasibility theory that applies to both MPC and RL by filling in the missing part of feasibility analysis for an arbitrary policy. The basis of our theory is to decouple policy solving and implementation into two temporal domains: virtual-time domain and real-time domain. This allows us to separately define initial and endless, state and policy feasibility, and their corresponding feasible regions. Based on these definitions, we analyze the containment relationships between different feasible regions, which enables us to describe the feasible region of an arbitrary policy. We further provide virtual-time constraint design rules along with a practical design tool called feasibility function that helps to achieve the maximum feasible region. The feasibility function either represents a control invariant set or aggregates infinite steps of constraints into a single one. We review most of existing constraint formulations and point out that they are essentially applications of feasibility functions in different forms. We demonstrate our feasibility theory by visualizing different feasible regions under both MPC and RL policies in an emergency braking control task.
keywords:
feasibility , constrained optimal control , model predictive control , reinforcement learning
@@
[aff1] organization=School of Vehicle and Mobility and State Key Lab of Intelligent Green Vehicle and Mobility, Tsinghua University, city=Beijing, postcode=100084, country=China,
@@
[aff2] organization=Department of Mechanical Engineering, University of California, Berkeley, postcode=94720, state=CA, country=USA,
@@
[aff3] organization=Robotics Institute, Carnegie Mellon University, city=Pittsburgh, postcode=15289, state=PA, country=USA,
1 Introduction
@@
Optimal control is an important theoretical framework for sequential decision-making and control. The goal of solving an optimal control problem (OCP) is to find a policy that maximizes some performance index, usually measured through cumulative rewards. In many real-world control tasks, such as robotics (Brunke et al., 2022), aerospace engineering (Ravaioli et al., 2022), and autonomous driving (Guan et al., 2022), the policy not only needs to optimize performance but also must take safety into consideration. Such problems can be formulated as constrained OCPs, where some state constraints must be strictly satisfied at every time step.
@@
In a constrained OCP, satisfying constraints at a single time step is not enough to ensure long-term safety. The policy may still run into a state where no constraint-satisfying solution can be found in future steps. This issue is called the infeasibility phenomenon (Li, 2023), which is one of the most important problems in constrained OCPs. Due to the existence of this phenomenon, we must ensure that any future state encountered by the policy is always feasible. In other words, the policy must work in a feasible region where the infeasibility phenomenon never happens. Therefore, it is necessary to identify the feasible region before implementing a policy in real-world environments. To rigorously define and analyze feasible regions, a set of feasibility theories for constrained OCPs is needed.
@@
Existing feasibility theories are built for model predictive control (MPC), which solves the optimal action online through receding horizon optimization. A central concept in these theories is persistent feasibility (Zhang et al., 2016; Borrelli et al., 2017), also referred to as recursive feasibility in some literature (Löfberg, 2012; Boccia et al., 2014), which describes long-term constraint satisfaction in a receding horizon control (RHC) problem. Specifically, an RHC problem is persistently feasible if its initially feasible set equals its maximal positive invariant set. Here, the initially feasible set contains all states where a constraint-satisfying solution can be found. The maximal positive invariant set contains all states whose successive trajectories always stay in this set. Persistent feasibility ensures that as long as there is a constraint-satisfying solution at the initial step, the receding horizon optimization can proceed without encountering the infeasibility phenomenon. A limitation of persistent feasibility is that it only applies to the optimal policy of a constrained OCP. This is because the maximal positive invariant set is defined in the closed-loop system under the MPC controller, which is the optimal controller of RHC problem. However, reinforcement learning (RL), as another important control method, solves the optimal policy in an iterative manner, which comes with a series of non-optimal intermediate policies. For such non-optimal policies, the infeasibility phenomenon may still be encountered even if the RHC problem is persistently feasible. It is necessary to analyze feasible regions of these non-optimal policies for improving safety through iteration, which is not available with the concept of persistent feasibility. Another concept called strong feasibility (Gondhalekar et al., 2009; Gondhalekar and Jones, 2011) expands the scope of persistent feasibility from the optimal policy to any initially feasible policy. An RHC problem is strongly feasible if, from every state in the initially feasible set, the trajectory under any feasible solution remains in this set. In other words, the feasible region of any initially feasible policy exactly equals the initially feasible set in a strongly feasible RHC problem. However, fulfilling strong feasibility requires a specific type of constraint design where the endlessly feasible region exactly equals the initially feasible region, a condition that may not be met for a general constrained OCP. Therefore, a more general set of theoretical tools is needed for analyzing the feasibility of an arbitrary policy.
@@
Figure 1: Illustration of core concepts in feasibility theory. The first two rows show state trajectories and feasible regions of an RL policy trained for 50 steps under a Hamilton-Jacobian (HJ) reachability constraint. The third row shows the results of 10000 training steps. The red marks represent endlessly feasible states, the blue marks for initially feasible states, and the gray marks for infeasible states. The purple-shaded areas represent real-time constraints, and the orange-shaded areas represent virtual-time constraints. A state with a red cross indicates violation of real-time or virtual-time constraints. The data in the figure comes from a numerical example of emergency braking control, and details can be found in Section 8.
@@
In order to fill this gap, we propose a set of feasibility theories for constrained OCPs that apply to both MPC and RL. Our theory is based on an extension of the temporal domain decoupling view in MPC to RL. We point out that any optimal control method follows the practice of solving OCP in virtual-time domain and implementing optimal policy in real-time domain. This decoupling view allows us to define the feasibility of an arbitrary policy, which is necessary for analyzing the feasible regions of intermediate non-optimal policies during RL training. Figure 1 illustrates the core concepts of our feasibility theory. Each real-time step corresponds to a virtual-time constrained OCP, whose optimal solution is either computed at each step online (like MPC) or solved offline before policy implementation (like RL). The constraint in virtual-time domain can be different from that in real-time domain. Whether the virtual-time constraint can be satisfied and how long it can be satisfied leads to different types of feasibility and feasible regions. For a typical safe RL algorithm, the feasible region of its policy starts as a small set and gradually expands during training, finally approaching the feasible region of the optimal policy. MPC, on the other hand, directly computes the optimal control sequences, and therefore, its feasible region is always the same. The main contributions of this paper are summarized as follows.
  • 1.
@@
    We propose a set of feasibility theories for constrained OCPs that apply to both MPC and RL. Based on a decoupling view of virtual-time and real-time domains, we separately define initial and endless feasibility, which describe short-term and long-term constraint satisfaction respectively. We further distinguish state and policy feasibility, which are defined based on the existence of a policy and on a given policy. By combining different types of feasibility, we derive four kinds of feasible regions. We further define the maximum feasible region and reveal its relationship to real-time constraints.
  • 2.
@@
    We give and prove containment relationships of all kinds of feasible regions, which provides a tool for analyzing the feasibility of an arbitrary policy. In particular, we analyze the containment relationship between a policy-specific feasible region and the maximum feasible region, which is a priority concern when solving a constrained OCP. For an arbitrary policy, its feasible region is also bounded from above and below, appearing as an intermediate between some known feasible regions.
  • 3.
@@
    We provide a collection of virtual-time constraint design rules along with a practical design tool called feasibility function that helps to achieve the maximum feasible region. Specifically, it is revealed by the above-mentioned containment relationships that the maximum feasible region of the optimal policy can be achieved by simply designing a maximum initially feasible region when initially and endlessly feasible regions are equal. The feasibility function ensures this equality by either constructing a control invariant set or aggregating infinite steps of constraints into a single-step one. We review most of existing constraint formulations and point out that they are essentially applications of feasibility functions in different forms.
@@
The rest of this paper is organized as follows. Section 2 introduces a decoupling view of two temporal domains. Section 3 illustrates the infeasibility phenomenon. We define feasibility and feasible regions in Section 4 and further analyze containment relationships of feasible regions in Section 5. After that, we provide feasibility function for constraint design in Section 6 and review existing constraint formulations in Section 7. Finally, Section 8 demonstrate feasible regions with experiments and Section 9 concludes the paper.
2 Constrained Optimal Control Problems in Two Temporal Domains
@@ 2.1 Constrained optimal control problems
$h(x_{t+i})\leq 0,i=0,1,2,\dots,\infty,$ (1)
where $h:\mathcal{X}\to\mathbb{R}$ is the constraint function. The constrained set is defined as $\mathrm{X}_{\mathrm{cstr}}=\{x\in\mathcal{X}|h(x)\leq 0\}$. Its complement is the unconstrained set, $\bar{\mathrm{X}}_{\mathrm{cstr}}=\mathcal{X}\setminus\mathrm{X}_{\mathrm{cstr}}$, which contains all states violating the constraints. Our aim is to find a policy $\pi:\mathcal{X}\to\mathcal{U}$ that maximizes the expected cumulative rewards under the state constraints,
$\max_{\pi}$ $\mathbb{E}_{x_{t}\sim d_{\mathrm{init}}(x)}\left\{\sum_{i=0}^{\infty}\gamma^{i}r(x_{t+i},u_{t+i})\right\},$ (2)
@@ 2.2 Real-time domain and virtual-time domain
To explain what feasibility is, it is necessary to recall the working mechanism of optimal control. In essence, any optimal controller works in two separated temporal domains, i.e., virtual-time domain and real-time domain. Under this perspective, the optimal policy is implemented in the real-time domain, while the solving algorithm is designed in the virtual-time domain. That is to say, constrained MPC/RL algorithm design should be discussed in the virtual-time domain, rather than in the real-time domain. This new viewpoint is our basis for understanding what feasibility is. The separation of the virtual-time domain from the real-time domain inspire us that a different constraint can be used in virtual-time domain from that in real-time domain since constrained OCP is not defined in the real-time domain. Therefore, one can choose a new constraint function in the virtual-time domain:
$g(x_{i|t})\leq 0,i=0,1,2,\dots,n,$ (3)
where $g(\cdot)$ is the virtual-time constraint function, and $n$ is the horizon length of virtual-time constrain. This length can be either finite or infinite. With virtual-time constraint (3), we can define the virtual-time OCP as follows.
$\max_{\pi}$ $\mathbb{E}_{x_{0|t}\sim d_{\mathrm{init}}(x)}\left\{\sum_{i=0}^{N}\gamma^{i}r(x_{i|t},u_{i|t})\right\},$ (4)
@@
where $N$ is the horizon length of virtual-time objective, which can be either finite (as in MPC) or infinite (as in RL).
It must be noted that (2) and (4) are not defined in the same temporal domain. The former is defined in the real-time domain, and the latter is defined in the virtual-time domain. The two domains are distinguished by their subscripts, of which $t+i$ represents the $(t+i)$-th step in the real-time domain and $i|t$ represents the $i$-th point in the virtual-time domain starting from time $t$. The virtual-time constraint (3) can be different from the real-time constraint (1) in two aspects: (1) constraint function and (2) time horizon. The real-time constraint is determined by the control task itself and cannot be modified by algorithm designers. Moreover, this constraint should be satisfied in infinite horizon because any real-world control task is continuing without termination. In contrast, the virtual-time constraint does not need to be posed at every step. It even does not need to have the same function form at all steps. That is to say, except at some special time instances, the constraints in virtual-time domain can be changed or even removed. In fact, a virtual-time constraint can be designed freely as long as it satisfies one requirement: it must be not weaker than the real-time constraint in the current step, i.e., if $g(x_{0|t})\leq 0$, then $h(x_{t})\leq 0$. Besides building the basic structure of defining feasibility, this domain-separating perspective provides us with great flexibility to build various virtual-time constraints, as well as the formulation of their constrained OCPs. As an example, a commonly used constraint design in MPC is that the first $n-1$ virtual-time constraints are the same as the real-time constraint $h$, and the last constraint $g(x_{n|t})$, called the terminal constraint, is designed differently, typically stronger than $h$ to ensure that $h$ is satisfied at every real-time step.
3 Illustration of Infeasibility Phenomenon
Infeasibility is the most important concept in constrained OCPs. The domain separation perspective can help us understand the infeasibility phenomenon more straightforwardly. Formally, “infeasibility” describes the phenomenon that the constraint of a virtual-time OCP cannot be satisfied by any policy or by a given policy at some state. Note that infeasibility concerns virtual-time OCP instead of real-time OCP because the policy is obtained from solving the former instead of the latter. We say a state is infeasible when no solution to the virtual-time OCP starting from it can be found. We say a policy is infeasible when it is not a valid solution to a virtual-time OCP. There are two reasons that lead to the infeasibility phenomenon: 1) insufficient training of policy in RL, and 2) improper design of virtual-time constraints.
@@ 3 Illustration of Infeasibility Phenomenon
To better explain how the infeasibility phenomenon occurs, let us consider the following example. Figure 1 shows an infeasible emergency braking control of an RL policy under a 1-step Hamilton-Jacobian (HJ) reachability constraint. This constraint requires the state not to enter the orange-shaded region in virtual-time domain. Here, the environment model is assumed to be perfect, i.e., the action and state trajectories are the same in the virtual-time and real-time domains. In the first two rows, at time $t=0$, the action given by the policy $\pi_{1}$ ensures that the virtual-time constraint is satisfied, i.e., $x_{1|0}\in\mathrm{X}_{\mathrm{cstr}}$. Since the model is perfect, this action transfers the current state to a new state $x_{1}$ in the real-time domain, which equals $x_{1|0}$. At time $t=1$, the action given by $\pi_{1}$ cannot satisfy the virtual-time constraint, $x_{1|1}\notin\mathrm{X}_{\mathrm{cstr}}$. This is how the infeasibility phenomenon occurs, and in this situation, we say $\pi_{1}$ is infeasible in state $x_{1}$. Note that constraint violation in virtual-time domain does not mean that there is no admissible action in real-time domain, where “admissible” means leading to a constraint-satisfying state. Actually, at time $t=1$, the action is still admissible in real-time domain because the resulting next state $x_{2}$ is still inside the real-time constrained set. However, since the virtual-time constraint is already violated, the violation of the real-time constraint is inevitable sometime in the future as long as the virtual-time constraint is properly designed. As shown in the figure, at time $t=N$, the next state $x_{N+1}$ finally goes out of the real-time constrained set.
@@
The above infeasibility phenomenon is mainly caused by insufficient training. Actually, extending the training step from 50 to 10000 basically solves the infeasibility problem. As shown in the third row in Figure 1, the state trajectory is fully contained in the virtual-time constrained set for all times, achieving a desired property called endless feasibility. The other reason that leads to the infeasibility phenomenon, i.e., improper design of virtual-time constraints, is illustrated in Figure 2. Here, the virtual-time constraint function is the same as the real-time one, i.e., $g=h$, but with a finite horizon $n$. This constraint is called a pointwise constraint. The policy is an MPC controller, which is the optimal policy of the virtual-time OCP. It shows that when $n$ is too small, infeasibility phenomenon occurs easily, and when $n$ is large enough, the state is always feasible. Usually, weaker and fewer steps of constraints have a worse ability to guarantee feasibility in the long run because they cannot provide sufficient confinement to future state trajectories. This problem cannot be mitigated by increasing training steps because even the optimal policy is infeasible. Regarding virtual-time constraint design, we provide a collection of constraint design rules along with a practical design tool called feasibility function that can avoid the infeasibility phenomenon and help to achieve the maximum feasible region.

### v2

@@ Abstract
Satisfying safety constraints is a priority concern when solving optimal control problems (OCPs). Due to the existence of infeasibility phenomenon, where a constraint-satisfying solution cannot be found, it is necessary to identify a feasible region before implementing a policy. Existing feasibility theories in model predictive control (MPC) only apply to the case where the policy is a solution to a specific OCP, either optimal or suboptimal. Such feasibility is essentially feasibility of OCPs, not that of policies or states. However, reinforcement learning (RL), as another important control method, represents a policy as a mapping from state to action, which itself is decoupled with OCPs. An RL policy may be not a solution to a specific OCP, i.e., not satisfying the OCP’s constraints, especially at an early stage of training. Feasibility analysis of such inadequately trained policies is necessary for safety improvement in RL; but that is not available under existing MPC feasibility theories. This paper proposes a feasibility theory that applies to both MPC and RL by decoupling states, constraints, and policies. Starting from a state, different constraints can be constructed, and different policies can be applied. We reveal that feasibility depends on the combination of these three elements, extending the traditional viewpoint of OCP-specific feasibility theories in MPC. The basis of our theory is to distinguish initial and endless, state and policy feasibility, and their corresponding feasible regions. Based on these concepts, we analyze the containment relationships between different feasible regions, which enables us to describe feasibility under arbitrary combinations of states, constraints, and policies. We further provide virtual-time constraint design rules along with a practical design tool called feasibility function that helps to achieve the maximum feasible region. The feasibility function either represents a control invariant set or aggregates infinite steps of constraints into a single one. We review most of existing constraint formulations and point out that they are essentially applications of feasibility functions in different forms. We demonstrate our feasibility theory by visualizing different feasible regions under both MPC and RL policies in an emergency braking control task.
@@
Yujie Yang
School of Vehicle and Mobility, Tsinghua University
yangyj21@mails.tsinghua.edu.cn and Zhilong Zheng
School of Vehicle and Mobility, Tsinghua University
zheng-zl22@mails.tsinghua.edu.cn and Masayoshi Tomizuka
Department of Mechanical Engineering,
University of California, Berkeley
tomizuka@berkeley.edu and Changliu Liu
Robotics Institute, Carnegie Mellon University
cliu6@andrew.cmu.edu and Shengbo Eben Li
School of Vehicle and Mobility & College of AI, Tsinghua University
@@
Y. Yang and Z. Zheng contributed equally to this work. All correspondence should be sent to S. E. Li.
@@
Chapter 1 Introduction
@@
Optimal control is an important theoretical framework for sequential decision-making and control. The goal of solving an optimal control problem (OCP) is to find a policy that maximizes some performance index, usually measured through cumulative rewards. In many real-world control tasks, such as robotics [brunke2022safe], aerospace engineering [ravaioli2022safe], and autonomous driving [guan2022integrated], policies must optimize performance while also guaranteeing safety. These problems are formulated as constrained OCPs, where certain state constraints must be strictly satisfied at every time step.
@@
In a constrained OCP, satisfying constraints at a single time step is insufficient to ensure long-term safety. A policy may still reach a state where no future constraint-satisfying actions exist. This issue is called the infeasibility phenomenon [li2023reinforcement]. To avoid this problem, the feasible region of a policy, where there is always a constraint-satisfying solution, must be identified before real-world deployment. Rigorously defining and analyzing feasible regions requires a general theory of feasibility for constrained OCPs.
@@
Existing feasibility theories are primarily developed for model predictive control (MPC), which computes optimal actions online through receding horizon optimization. A central concept in these theories is persistent feasibility [zhang2016switched, borrelli2017predictive], also known as recursive feasibility [lofberg2012oops, boccia2014stability], which describes long-term constraint satisfaction in a receding horizon control (RHC) problem. Specifically, an RHC problem is persistently feasible if its initially feasible set equals its maximal positive invariant set. Under this condition, any initial constraint-satisfying solution ensures that subsequent receding horizon optimizations will not encounter infeasibility. Persistent feasibility only applies to the optimal policy of a constrained OCP because its maximal positive invariant set is defined for the closed-loop system under the optimal MPC controller. The concept of strong feasibility [gondhalekar2009controlled, gondhalekar2011mpc] and suboptimal MPC [scokaert1999suboptimal, pannocchia2011conditions] expand persistent feasibility from the optimal policy to any initially feasible policy. They guarantee long-term safety not by requiring an optimal solution, but merely a feasible solution to the RHC problem at each step. Specifically, an RHC problem is strongly feasible if, from every state in the initially feasible set, the trajectory under any feasible solution remains in this set.
A key limitation of MPC-based feasibility theories is that they only apply to policies that are solutions to a specific OCP—either optimal or suboptimal—meaning the policy inherently satisfies the OCP’s constraints. In this context, the policy is coupled with the OCP, and the feasibility analysis pertains to the problem itself, not to the policy or state in isolation. However, reinforcement learning (RL), as another important control method, represents a policy as a mapping from state to action, decoupling it from any single OCP. Although RL also solves an OCP, it does not obtain the optimal solution until training converges. During early training stages, an RL policy may not even be a feasible solution to the intended OCP. Feasibility analysis of such inadequately trained policies is crucial for safety improvement in RL, yet this falls outside the scope of existing MPC feasibility theories.
There are a few existing works on safe RL that mention the concept of feasibility and leverage it to guarantee long-term safety. For example, ma2021feasible define feasible states as those from which the cumulative cost constraints can be satisfied under a given policy, and introduce the feasible region as the set of all states that are feasible under any policy. yu2022reachability define the feasible set of a policy as the set of states where infinite-horizon state constraints can be satisfied, and introduce the largest feasible set as the union of feasible sets over all policies. [zheng2024safe] follow the definition of yu2022reachability and extend the application of feasibility to offline settings. However, these works have not yet formed a systematic feasibility framework and use varying, often inconsistent, terminology. More importantly, their notion of feasibility refers to the satisfaction of real-world safety constraints during policy execution, which differs from the classical MPC notion of feasibility, i.e., the existence of a solution to the OCP. In the context of RL, feasibility should also consider the existence of solutions to the underlying OCP during the training phase, where algorithmic constraints may differ from those in deployment. This aspect remains under-explored in existing safe RL literature and constitutes a major focus of our work.
@@
Figure 1.1: Illustration of core concepts in feasibility theory. The second row shows state trajectories and feasible regions of an RL policy trained for 50 steps under a Hamilton-Jacobi reachability constraint, with the first row demonstrating the first virtual-time step corresponding to each real-time step. The rest grey shaded squares are stacked along the virtual-time axis. The third row shows the results of 10000 training steps. The purple-shaded areas represent real-time constraints, and the orange-shaded areas represent virtual-time constraints. The circles stand for the states at every step on a trajectory. The squares and diamonds stand for regular states so as to show different feasibility regions. Regardless of the shape, all blue marks represent initially feasible states, all red marks for endlessly feasible states, and all gray marks for infeasible states. Intuitively, “initially feasible” means that there exists a policy satisfying the virtual-time constraint at the current step, while “endlessly feasible” means that such a policy will always exist. A state with a red cross indicates violation of real-time or virtual-time constraints. The data in the figure comes from a numerical example of emergency braking control, and details can be found in Chapter 8.
@@
In order to fill this gap, we propose a set of feasibility theories for constrained OCPs that apply to both MPC and RL. The key to our theory is a decoupling view of states, constraints, and policies in constrained OCPs. Starting from a state, different constraints can be constructed, and different policies can be applied. We reveal that feasibility depends on the combination of these three elements instead of just on the OCP itself, which extends the traditional viewpoint of OCP-specific feasibility theories in MPC. Figure 1.1 illustrates the core concepts of our feasibility theory. Any optimal control method follows the practice of solving an OCP in the virtual-time domain and implementing a policy in the real-time domain. Each real-time step corresponds to a virtual-time constrained OCP, solved either at each step online (as in MPC) or offline before deployment (as in RL). In the real-time domain, the control task naturally has some intrinsic constraints, termed real-time constraints, that are immutable by human design, such as collision avoidance. However, the original real-time constraints may be challenging to solve. As a result, in the virtual-time domain, it is typically necessary to design the virtual-time constraints differing from the real-time ones to facilitate solving. The satisfiability of virtual-time constraints over time defines different types of feasibility and feasible regions. For a typical safe RL algorithm, the feasible region of its policy starts as a small set and gradually expands during training, eventually approaching that of the optimal policy. This is because the policy is typically initialized with random parameters, leading to unpredictable and often unsafe actions, which corresponds to a feasible region. As training proceeds, the policy is explicitly optimized to maximize reward while minimizing constraint violations. This process gradually guides the policy towards safer behaviors, effectively expanding the feasible region. In contrast, MPC, which directly computes optimal control sequences, typically maintains a fixed feasible region. The main contributions of this paper are summarized as follows.
@@
  • We propose a set of feasibility theories for constrained OCPs that apply to both MPC and RL. By decoupling states, constraints, and policies, we define separate notions of initial and endless feasibility for both states and policies, characterizing short-term and long-term safety with and without a specific policy. Based on these, we derive four kinds of feasible regions and establish the relationship between the maximum feasible region and real-time constraints.
@@
  • We derive and prove containment relationships among all feasible regions, which provides a tool for analyzing feasibility for arbitrary combinations of states, constraints, and policies. In particular, we analyze the relationship between a policy-specific feasible region and the maximum feasible region, which is a primary concern when solving a constrained OCP. We also show that for any policy, its feasible region is bounded above and below by known feasible regions.
@@
  • We provide practical rules for designing virtual-time constraints and introduce a tool called feasibility function to help achieve the maximum feasible region. Specifically, the maximum feasible region is achieved by designing a maximum initially feasible region when initially and endlessly feasible regions are equal. The feasibility function guarantees this equality by either constructing a control invariant set or aggregating infinite-horizon constraints into a single step. We review existing constraint formulations and show that they are essentially specific applications of feasibility functions.
It is worth emphasizing that what we intend to propose in this paper is a feasibility theory regarding how different designs of virtual-time constraint influence the feasibility property of the solution of an constrained OCP, rather than a specific solving method. Hence, this paper does not cover discussions on how to obtain or to what extent we can obtain the optimal solution of an constrained OCP.
@@
The rest of this paper is organized as follows. Chapter 2 introduces a decoupling view of two temporal domains. Chapter 3 illustrates the infeasibility phenomenon. We define feasibility and feasible regions in Chapter 4 and analyze their containment relationships in Chapter 5, we then introduce the feasibility function for constraint design in Chapter 6 and review existing constraint formulations in Chapter 7. Finally, Chapter 8 demonstrates feasible regions with experiments, and Chapter 9 concludes the paper.
Chapter 2 Constrained Optimal Control Problems in Two Temporal Domains
@@ 2.1 Constrained optimal control problems
$h(x_{t+i})\leq 0,i=0,1,2,\dots,\infty,$ (2.1)
where $h:\mathcal{X}\to\mathbb{R}$ is the constraint function. The constrained set is defined as $\mathrm{X}_{\mathrm{cstr}}=\{x\in\mathcal{X}|h(x)\leq 0\}$. Its complement, $\bar{\mathrm{X}}_{\mathrm{cstr}}=\mathcal{X}\setminus\mathrm{X}_{\mathrm{cstr}}$, contains all states violating the constraints. Our aim is to find a policy $\pi:\mathcal{X}\to\mathcal{U}$ that maximizes the expected cumulative rewards under the state constraints,
$\max_{\pi}$ $\mathbb{E}_{x_{t}\sim d_{\mathrm{init}}(x)}\left\{\sum_{i=0}^{\infty}\gamma^{i}r(x_{t+i},u_{t+i})\right\},$ (2.2)
@@ 2.2 Real-time domain and virtual-time domain
To explain feasibility, it is necessary to recall the working mechanism of optimal control. In essence, any optimal controller works in two separated temporal domains: virtual-time domain and real-time domain. In virtual-time domain, an OCP is solved to obtain an optimal policy, and in real-time domain, the optimal policy is implemented. Virtual-time domain and real-time domain have their own OCP. The above OCP (2.2) is actually defined in real-time domain. The motivation for defining another OCP in virtual-time domain is that OCP (2.2) is difficult to solve due to its infinite-step constraint. Introducing virtual-time OCP allows us to replace this constraint with a finite-step one:
$g(x_{i|t})\leq 0,i=0,1,2,\dots,n,$ (2.3)
where $g(\cdot)$ is the virtual-time constraint function, and $n$ is its finite horizon length. With virtual-time constraint (2.3), we can define the virtual-time OCP as follows.
$\max_{\pi}$ $\mathbb{E}_{x_{0|t}\sim d_{\mathrm{init}}(x)}\left\{\sum_{i=0}^{N}\gamma^{i}r(x_{i|t},u_{i|t})\right\},$ (2.4)
@@
where $N$ is the horizon length of virtual-time objective function, which can be either finite (as in MPC) or infinite (as in RL).
It must be noted that OCP (2.2) is defined in the real-time domain, and OCP (2.4) is defined in the virtual-time domain. The two domains are distinguished by their subscripts: $t+i$ represents the $(t+i)$-th step in the real-time domain, and $i|t$ represents the $i$-th point in the virtual-time domain starting from time $t$. The virtual-time constraint (2.3) can be different from the real-time constraint (2.1) in two aspects: 1) constraint function and 2) time horizon. The real-time constraint is determined by the control task itself and cannot be modified by algorithm designers. Moreover, it always has an infinite horizon because any real-world control task is continuous without termination. In contrast, the virtual-time constraint is constructed by algorithm designers and has a finite horizon $n$, which is different from the horizon length of virtual-time objective function $N$. There is only one requirement for virtual-time constraint: it must be not weaker than the real-time constraint in the current step, i.e., if $g(x_{0|t})\leq 0$, then $h(x_{t})\leq 0$.
Recall the commonly used constrained optimal control formulation in MPC:
$\min_{u_{0|t},u_{1|t},\dots,u_{N|t}}$ $\sum_{i=0}^{N-1}l(x_{i|t},u_{i|t})+l_{T}(x_{N|t},u_{N|t}),$ (2.5)
s.t. $x_{i+1|t}=f(x_{i|t},u_{i|t}),$
$h(x_{i|t})\leq 0,i=0,1,2,\dots,N-1,$
$g_{T}(x_{N|t})\leq 0,$
where $l(\cdot,\cdot)$ is the utility function. Problem (2.5) is actually a variant of the virtual-time OCP (2.4). Specifically, 1) the horizon length of the constraint function in problem (2.5) equals that of the objective function, i.e., $n=N$, 2) the maximization of rewards is changed to the minimization of cost, 3) the first $n$ steps of the virtual-time constraints equals the real-time constraints, i.e., $g(x_{i|t})=h(x_{t+i}),i=0,1,2,\dots,N-1$, 4) the last utility is a specially designed terminal cost $l_{T}(\cdot)$, and 5) the last constraint is a specially designed terminal constraint $g_{T}(\cdot)$, typically stronger than $h$ to ensure that $h$ is satisfied at every real-time step.
We visualize the trajectories of an MPC controller obtained by solving problem (2.5) in an emergency braking task in Figure 2.1. Details of the task can be found in Chapter 8. The prediction horizon is 5 and there is no terminal constraint. It can be observed that the two trajectories always coincide at the first two steps since the first virtual-time action is implemented in the real-time domain. Beyond that, the two trajectories diverge when the virtual-time trajectory starts from $t=2$ and $t=3$. This is because the prediction horizon is not long enough to take into account all future constraints, leading to a short-sighted objective-seeking policy. The virtual-time trajectory starting from $t=4$ and $t=5$ overlap with the real-time trajectory because in both cases the vehicle is doing best-effort control to avoid constraint violation, which is already impossible given the current state and control limits. This illustrates the fundamental difference between virtual-time and real-time domains in constrained OCPs.
Figure 2.1: Real-time and virtual-time trajectories of an MPC controller in an emergency braking task. The real-time trajectories are the same across the four figures. The virtual-time trajectories starts at different real-time steps. The gray-shaded area represents the constraint-violating set. The red-shaded area represents the set where there exists a policy to ensure infinite-horizon safety. A state with a red cross indicates violation of the real-time constraint.
Chapter 3 Illustration of the Infeasibility Phenomenon
Infeasibility is the most important concept in constrained OCPs. It describes the phenomenon that the constraint of a virtual-time OCP cannot be satisfied. Formally, we define infeasibility as follows.
Definition 3.0.1 (Infeasibility).
Consider a virtual-time OCP with virtual-time constraint $g(\cdot)\leq 0$ over horizon $i=0,\dots,n$.
  1. 1.
    A state $x$ is infeasible (w.r.t. $g$) if there exists no policy $\pi$ that is a solution to the virtual-time OCP.
  2. 2.
    A policy $\pi$ is infeasible in state $x$ (w.r.t. $g$) if it is not a solution to the virtual-time OCP.
There are two reasons for the infeasibility phenomenon: 1) imperfect policy solving and 2) improper constraint design.
@@
To better explain how the infeasibility phenomenon occurs, let us consider the following example. Figure 1.1 shows an infeasible emergency braking control of an RL policy under a one-step constraint that requires the state not to enter the orange region. In the first two rows, at $t=0$, policy $\pi_{1}$ satisfies the virtual-time constraint, i.e., $x_{1|0}$ is not in the orange region. The state is transferred to $x_{1}$ in the real-time domain, which equals $x_{1|0}$. At $t=1$, $\pi_{1}$ cannot satisfy the virtual-time constraint, i.e., $x_{1|1}$ enters the orange region. This is how the infeasibility phenomenon occurs. In this situation, we say $\pi_{1}$ is infeasible in state $x_{1}$. Note that constraint violation in virtual-time domain does not mean that there is no admissible action in real-time domain, where “admissible” means leading to a constraint-satisfying state. At $t=1$, the action is still admissible in real-time domain because the resulting next state $x_{2}$ is still inside the real-time constrained set. However, since the virtual-time constraint is already violated, the violation of the real-time constraint is inevitable sometime in the future. As shown in the figure, at time $t=N$, the next state $x_{N+1}$ finally goes out of the real-time constrained set.
@@
The above infeasibility phenomenon is mainly caused by imperfect policy solving. Extending the training step from 50 to 10000 basically solves the infeasibility problem. As shown in the third row in Figure 1.1, the state trajectory is entirely contained in the virtual-time constrained set. More detailed examples can be found in Chapter 8, where we demonstrate through an emergency braking task that the feasible regions of RL policies expand as training proceeds, indicating that higher solution quality leads to better feasibility. Another reason for the infeasibility phenomenon, i.e., improper constraint design, is illustrated in Figure 3.1. Here, the virtual-time constraint function is a pointwise constraint, i.e., $g=h$ with a finite horizon $n$. The policy is an MPC controller. It shows that the infeasibility phenomenon occurs when $n$ is too small and disappears when $n$ is large enough. More detailed examples on the effect of constraint design can also be found in Chapter 8, where we show that with an MPC controller, different virtual-time constraints leads to different feasible regions. Intuitively, weaker and fewer steps of virtual-time constraints have worse feasibility in the long run because they cannot provide sufficient confinement to future state trajectories. This intuition is formally stated and proved in Section 5.1.
Regarding virtual-time constraint design, we provide a collection of design rules and a practical design tool called feasibility function in Chapter 6. This tool helps avoid the infeasibility phenomenon and achieve the maximum feasible region.